OpenStudy (anonymous):

find the exact values of theta cos theta=-1/4 tan theta >0 cot theta= 4/3 cos theta <0 show me step by step how to do this

OpenStudy (campbell_st):

cos is -ve and tan +ve.. the problem is in the 3rd quadrant... 3rd quadrant the angle is \[180 + \theta\] find

OpenStudy (campbell_st):

find cos ^(-1) (1/4) = and add to 180

OpenStudy (campbell_st):

cot = 4/3 cot = 1/tan tan +ve and cos -ve means 3rd quadrant again find tan^-1(3/4) and then add the angle to 180